# Shear flow of frictional spheroids: lentil-like grains slide more than rice-like ones

### v1

@@ I Introduction
To capture the physics of these shape-dependent behaviors, various modeling approaches have been developed. The classical $\mu(I)$ rheology, originally formulated for spheres [24], has been extended to frictionless spheroids [25], to account for kinematic orientation [26]. The effect of shape has also been incorporated in the kinetic theory of grains [27]. These models serve as important tools for linking microscopic particle interactions to macroscopic flow properties, but they still lack applicability to dense flows of frictional particles.
Phase diagrams describing dissipation mechanisms as a function of particle inertia and friction have evolved from early studies on 2D granular spheres [28] to more complex dense suspensions [29] and 3D spherocylinders [22], which showed an extended sliding regime. Yet, the regime diagram for more complex shapes is missing, and predicting it does not appear trivial.
@@ II Numerical Protocol
${\mathcal{X}^{2}+\mathcal{Y}^{2}\over a^{2}}+{\mathcal{Z}^{2}\over c^{2}}=1\,,$ (1)
@@
where $\Omega$ is the volume of the system, $N$ is the number of particles, $\mathbf{r}^{(k)}$ the particle position vector and $\mathbf{F}^{(lk)}$ the force acting on particle $l$ due to the interaction with particle $k$. We neglect the contribution of the velocity fluctuations to the stress, since in the dense flows we are investigating (packing fraction $\phi>0.5$) it is at least two orders of magnitude smaller than the collisional one [45], which we also confirmed by computing it in post-processing.
We control the normal pressure component $\sigma_{yy}=-P$, by assigning a velocity $dL_{y}/dt=k_{p}(P_{t}-\overline{P})$ to the box edges with normals ±$\pm\mathbf{e}_{y}$, where ${\overline{P}}$ is the average pressure calculated between the control inputs, $P_{t}$ is the target vertical pressure, and $k_{p}$ is the proportional control gain. We set the target pressure so that the contact stiffness number $\kappa=\min(\alpha,1)(E/P)^{2/3}>1.1\cdot 10^{3}$, where $E$ is the Young modulus of the grains. Compared to [19] we scale the contact stiffness number by the aspect ratio, in order to guarantee a small overlap even for oblate particles. Doing so ensures we are in the rigid grain limit [19, 46] and the relevant dimensionless number is the inertial number $I=\dot{\gamma}d_{eq}\sqrt{\rho_{p}/P}$, where $\dot{\gamma}$ is the shear rate, $\rho_{p}$ is the particle density, $d_{eq}=2a\alpha^{1/3}$ is the equivalent volume sphere diameter. We set the Poisson’s ratio $\nu=0.3$, which sets the ratio of tangential to normal stiffness $k_{t}/k_{n}\approx 0.41$, see Appendix A for more detail.
@@
$N=2000$ particles are inserted in the box, with approximately $12$ particles per dimension and, with a $20\%$ degree of polydispersity for the value of the smallest semi-axis around the mean and none in the aspect ratio. All simulations start at an initial concentration $\phi_{0}=0.4$ with isotropic random distribution and orientation of grains, equal edges of the box $L_{x}=L_{y}=L_{z}$, and then the box is compressed in the $y$ direction to reach the desired pressure. The chosen number of particles is sufficient to characterize the rheology, see Appendix C. We set the coefficient of restitution $e=0.1$, insofar in the dense flow regime, the influence of $e$ on the effective friction of the system is found negligible [19], even though it could affect the particle-level dissipation [47].
@@ III.1 Alignment and Rotation
[…]
@@
where $\mathbf{I}$ is the identity matrix, and $\psi(\mathbf{p})$ is the probability distribution of $\mathbf{p}$ on the unit sphere, $\mathbb{S}^{2}$. The uniaxial nematic order parameter is defined as $S_{2}=3\lambda_{\text{max}}/2$, where $\lambda_{\text{max}}$ is the maximum eigenvalue of $\mathbf{Q}$. As shown in Figure 2a, $S_{2}$ increases from 00 for an isotropic distribution when $r_{g}\to 0$, to values close to $1$ for elongated and flattened particles, where a value $S_{2}=1$ means that all particles are aligned in the same direction. Additionally, we define as weakly anisotropic configurations where $0.350<S_{2}<$ 0.35.
@@
One might expect that microscopic friction broadens the distribution of particle orientations, thereby reducing nematic order. However, for particles with low to moderate eccentricity, the highest degree of nematic ordering is not observed in the frictionless limit, but at an intermediate friction coefficient of $\mu_{p}=0.1$. This underscores the dual role of friction: while it typically disrupts alignment in the nematic phase, it can promote alignment in the isotropic phase via torque-induced reorientation. Finally, notice how for small eccentricity, $r_{g}<$ the order parameter is within the weakly anisotropic case. Interestingly, oblate particles align slightly better up to moderate values of friction coefficient $\mu_{p}\leq 0.4$, but worse at high friction values with respect to their reciprocal prolate ones.
@@
As reported in several studies [14, 8, 20, 21, 22, 18, 10] grains do not simply align in the direction of least resistance to the flow, along the streamlines, but at a certain angle in the extension direction. Defining the mean angle of inclination of the major axis with respect to the flow direction projected in the $xy$ plane, $\theta_{x}$, 111for oblate ellipsoids since there are are two major axes we track the symmetry axis and then compute the corresponding angle for the major axis as $\pi/2-\theta_{x}$, we report data in Figure 2b. The higher $|r_{g}|$ the more the particles orient their major axis parallel to the flow direction. The effect of friction is also highly non-linear: $\theta_{x}$ is lower for $100.1<\mu_{p}<$ 10 at small degrees of eccentricity, but for high flattening and elongation, the minimum $\theta_{x}$ is found for high values of friction, when alignment is less intense. The exact reason why this alignment occurs at a slight angle with the streamlines is still unclear, but recent progress links the preferential angle to the noise provided by collisions with other particles in the shear flow [49].
@@
Another way to look at the emergence of alignment is to track the the average vorticity-component of angular velocity $\langle\omega_{z}\rangle$, see Figure 2c. When friction is present, spherical particles rotate on average as a fluid particle in shear flow [19] $\omega_{z}\approx\dot{\gamma}/2$. For non-spherical particles, elongation provides a geometric constraint, effectively hampering their rotation. For all values of friction, oblate particle rotate slightly less than prolate ones for the same degree of eccentricity. One possible explanation is that while some prolate particles orient their major axis along the vorticity direction, resulting in a log-rolling behavior, this is not observed for oblate particles (data not shown). This behavior is analogous to that seen in simple shear flow in a dilute viscous fluid, where rotation about the vorticity axis is stable for prolate spheroids when their minor axis aligns with it, but for oblate spheroids, stability occurs when their major axis is aligned instead [50]. In the high friction limit, yellow curve, the rotation approaches the one of the corresponding fluid particle, regardless of the shape. This is due to the lower packing fraction attained in this scenario, see Figure 7. To conclude, for non-spherical we notice that the average rotation is anticorrelated with the orientational order [7], but not directly related to the particles’ preferred orientation.
@@ III.2 Dissipation Regime diagram and fluctuations
As a function of the inertial number and the microscopic friction, the leading dissipative mechanism of granular flows and dense suspensions can be either due to normal collisions or to frictional sliding [28, 29, 22].
We can split the dissipated power into a collisional term due to contact inelasticity, $\mathcal{P}_{n}$, and a tangential term due to frictional sliding $\mathcal{P}_{t}$, in formulas:
@@
where $F_{i}$ is the force due to contact $i$, $N_{c}$ is the total number of contacts, $u_{i}$ is the relative velocity of the grains at the contact point, $F^{n,\gamma}$ is the normal component of the force due to damping and $N_{s}$ is the number of sliding contacts, those which satisfy $F_{i}^{t}=\mu_{p}F_{i}^{n}$. Figure 3 shows for each aspect ratio the regime diagram of the dissipation as the level set curve where $\mathcal{P}_{t}=\mathcal{P}_{n}$. The parametric space has three regions: the collisional regime for low $\mu_{p}$ and high $I$, the collisional-frictional or rolling regime for high $\mu_{p}$ and high $I$ and the frictional sliding regime for intermediate $\mu_{p}$ and low enough $I$. For $\alpha>1$ the diagram closely resembles the one obtained for sphero-cylinders in [22]. Here, we extend this diagram by also including $\alpha<1$. For values of $|M_{\alpha}|>1.2$ the crossover from the collisional to the sliding regime occurs at higher inertial number the higher the absolute value of the shape ratio (see the inset). Regarding the boundary between the sliding and collisional-frictional regime, non-spherical particles start to roll at higher values of $I$ and $\mu_{p}$.
Figure 3: Regime diagram of granular shear flow as a level set curve where the frictional and collisional dissipation are equal. Left lines mark the transition between the collisional and the sliding frictional regime, while right ones the transition to the collisional-frictional regime. The black dashed line has slope 2 as predicted by [28] for spheres. For all non-spherical particles the sliding regime is wider than for spherical ones, indicating non-spherical particles slide over a larger range of parameters. (Inset) Slice of $I_{c}$ at the collisional-sliding transition for $\mu_{p}=0.01$, showing non monotonicity in $\alpha$. Slightly aspherical particles display a reduced collisional regime.
To rationalize our findings we visualize the power dissipation normalized by the injected power into the system, given by $\dot{\gamma}\sigma_{xy}\Omega\sim\dot{\gamma}P\Omega$, in Figure 4. We observe that, up to moderate friction coefficients ($\mu_{p}\leq 0.4$), Figure 4a, the power dissipated due to collisions decreases as particles become more elongated or flattened for aspect ratios $M_{\alpha}>1.2$.
Owing to the system’s structural anisotropy, assuming all particles are perfectly aligned with their major axis along the flow direction, the cross-sectional areas of each grain are listed in Table 1
@@
Oblate $\alpha$ $1$ $\alpha$ $\alpha$ $1$
Prolate $1$ $\alpha$ $\alpha$ $1/\alpha$ $1/\alpha$
Among them, the gradient-vorticity cross-section $A_{yz}$ is the smallest, implying that collisions are least likely in the flow direction. Defining the particles’ velocity fluctuation as the excess over the linear velocity profile $\delta\bm{v}=\bm{v}-\dot{\gamma}y\mathbf{e}_{x}$, velocity fluctuations along the flow direction, $\delta v_{x}$, tend to persist longer, due to the reduced cross-sectional area. Moreover, as reported in [16] at high volume fractions and seen in Figure 5a and b, the dominant fluctuation is $\delta v_{x}$, indicating that most of the fluctuating kinetic energy is in the flow direction. Taken together, these effects offer a simple and consistent explanation for the reduced collisional power dissipation observed in systems with anisotropic particles: lower collision frequency along the dominant fluctuation direction leads to less energy lost through collisions. This simple geometrical argument also clarifies why, for flattened particles, fluctuations in the vorticity direction are of the same magnitude as those in the flow direction, as their respective cross-sections are equal $A_{xz}=A_{xy}$.
@@
Previous scaling arguments for circular particles estimated that the cross-over from collisional to sliding regime occurs at $I_{c}\sim\mu_{p}^{2}$ [28], as indicated by the dashed line in Figure 3. In our case, we find roughly $I_{c}\sim f(\alpha)\mu_{p}^{2}$, where $f(\alpha)$ a non-monotonic function, with global minimum in $\alpha=1.0$. The reduced collisional dissipation of aligned particles, shifts the transition to higher $I_{c}$ for both elongated [22] and flattened particles. For the latter, their slightly higher alignment shifts the transition at slightly higher inertial number compared to the respective reciprocal shapes.
@@
From Figure 3, we also observe that at a fixed inertial number $I$, the transition from rolling to sliding regime occurs at higher values of $\mu_{p}$ for non spherical particles. Oblate shapes start rolling at higher $\mu_{p}$ than prolate ones with similar eccentricity, as they experience more anisotropic fabric and comparatively stronger forces (Sections III.5, III.6). Interestingly, this second transition from sliding to rolling is monotonic across all $M_{\alpha}$, indicating that in the collisional rolling regime, lower volume fraction and less collective motion reduce geometric constraints; see Sections III.3 and III.4.
@@
Figure 4: (a) Normalized normal and (b) tangential dissipation for $I\approx 10^{-2}$ and different values of microscopic friction. On the left, data at $\mu_{p}<0.4$, on the right $\mu_{p}\geq 0.4$. In panel (a.1), the red (black) line represents the best-fit power law for frictionless prolate (oblate) particles, with a slope of $-0.90$, $(0.96)$. For particles with $\mu_{p}=0.1$, panel (b.1), the corresponding slopes are $-0.30$ for prolate and $+0.33$ for oblate particles. From (b.1, b.2), it is evident that the tangential dissipation increases up to a value of $\mu_{p}$ between $0.1$ and $1.0$, depending on the shape, and then reduces for higher values. Same color legend as in Figure 2.
@@
Figure 5: (a) Ratio of the streamline-wise velocity fluctuation $\delta v_{x}$ to the gradient-wise velocity fluctuation $\delta v_{y}$, (b) and to spanwise vorticity $\delta v_{z}$. (c) Ratio of rotational to translational kinetic energy, where the dashed line indicates a ratio of one. Different microscopic friction coefficients, with the same color legend as in Figure 4. For low values of friction the flow component of the fluctuation dominates for prolate particles, whereas for oblate ones, both the flow and the vorticity ones are dominant. In the sliding regime the fluctuations in the vorticity direction are reduced, as friction stabilizes that translation. For low values of friction and eccentricity the rotational fluctuations dominate. At high values of friction energy tends to be equally distributed between translation and rotation, as well as between the translational degrees of freedom. Data taken at $I\approx 0.01$. Same color legend as in Figure 2.
@@
Figure 6: (a.1-6) Intensity of normalized velocity fluctuations, $\delta v=|\delta\bm{v}|$, and (b.1-6) angular velocity fluctuations, $\delta\omega=|\delta\bm{\omega}|$. Results are shown for $\mu_{p}=0$ (left), $0.4$ (center), and $10.0$ (right). Black lines indicate best-fit power laws for spherical particles, with slopes of $-0.49$, $-0.19$, and $-0.46$ (a.1-6), and $-0.54$, $-0.25$, and $-0.47$ (b.1-6), where in the frictionless case the fit is performed on $\alpha=1.5$. Splitting the data for elongated and flattened particles makes the symmetry of reciprocal shapes less evident, but clearly shows the monotonic trend.
Next, we turn our attention to the non-monotonic transition from collisional to sliding regime of particles of small elongation, i.e. $M_{\alpha}\leq 1.2$, in the collisional regime. Previous studies showed how even a small degree of eccentricity, dramatically changes the microscopic dynamics both in static and dynamic conditions [51, 52]. Firstly, we recall that in all the parametric space, they mostly reach a weakly nematic phase (Figure 2b). Therefore, for these grains the caging effect is smaller and their rotation is less frustrated. We measure the fluctuating kinetic energy, that we decompose in the translational component $\delta E_{transl}=0.5\sum_{i=1}^{N}m_{i}||\delta\bm{v}_{i}||^{2}$ and the rotational one $\delta E_{rot}=0.5\sum_{i=1}^{N}\delta\bm{\omega}_{i}\cdot\mathbf{J}_{i}\delta\bm{\omega}_{i}$, where $\mathbf{J}_{i}$ is the moment of inertia tensor of the $i$-th particle, and $\delta\bm{\omega}$ its fluctuating angular velocity, measured as $\delta\bm{\omega}_{i}=\bm{\omega}_{i}-\langle\bm{\omega}\rangle$. Figure 5c shows that for low friction, $\mu_{p}<0.01$, the fluctuating rotational energy exceeds the translational one. Therefore, the previous scaling arguments [28, 29] exclusively based on the translational fluctuating energy need to be revisited to determine $\mathcal{P}_{n}$ in the case of slightly aspherical particles. One does also notice how for highly eccentric particles the rotational component is considerably smaller for $\mu\leq 0.4$, which makes the scaling arguments in [28] still applicable even for non-spherical particles. This motivates why the collisional to sliding transition lines have approximately the same slope, recall Figure 3. Notably, for high microscopic friction the two components of the fluctuating kinetic energy tend to attain a similar magnitude, indicating a thermalizing effect, which is also reflected in the velocity fluctuations.
@@
In Figure 6 we report the intensity of velocity fluctuations in the different regimes. Defining $\mathcal{L}=\delta v/\dot{\gamma}d_{eq}$, the so-called lever effect, which measures the dimensionless intensity of the velocity fluctuations ($\delta v$), in the collisional and collisional-frictional regimes, $\mathcal{L}\sim I^{-\xi}$, with $\xi\approx 0.5$ as reported by multiple studies [46, 28, 29, 53, 54]. In the sliding regime however, we find $\mathcal{L}\sim I^{-0.19}$ for spherical particles, slightly lower than the one reported in 2D [28]. For elongated particles, the fluctuations are reduced even more in the sliding regime, confirming the more coherent flow they undergo. In the frictional-rolling regime the shape of the particles has very little influence on the velocity fluctuations, signaling that all particles behave in a similar way as alignment is reduced, and the fluctuations tend to become isotropic, recall Figure 5.

### v2

# Shear flow of frictional spheroids: Comparison between elongated and flattened particles

@@ I Introduction
To capture the physics of these shape-dependent behaviors, various modeling approaches have been developed. The classical $\mu(I)$ rheology, originally formulated for spheres [24], has been extended to frictionless spheroids [25], to account for kinematic orientation [26]. The effect of shape has also been incorporated in the kinetic theory of both frictionless [27, 28] and frictional grains [29, 30], of non-spherical particles via an effective coefficient of restitution to account for frictional dissipation and both order tensors and elastic stresses. These models provide a high degree of accuracy with only a couple of fitting parameters, yet their application relies on the measurements of critical volume fraction, which can be challenging to determine under constant stress conditions [31].
Regime diagrams describing dissipation mechanisms as a function of inertial number and friction have evolved from early studies on 2D granular disks [32] to more complex dense suspensions [33] and 3D spherocylinders [22], which showed an extended sliding regime. Yet, the regime diagram for more complex shapes is missing, and predicting it does not appear trivial.
@@ II Numerical Protocol
${\mathcal{X}^{2}+\mathcal{Y}^{2}\over a^{2}}+{\mathcal{Z}^{2}\over c^{2}}=1\,.$ (1)
@@
where $\Omega$ is the volume of the system, $N$ is the number of particles, $\mathbf{r}^{(k)}$ the particle position vector and $\mathbf{F}^{(lk)}$ the force acting on particle $l$ due to the interaction with particle $k$. We neglect the contribution of the velocity fluctuations to the stress, since in the dense flows we are investigating (packing fraction $\phi>0.5$) it is at least two orders of magnitude smaller than the contact stress [49], which we also confirmed by computing it in post-processing.
We control the normal pressure component $\sigma_{yy}=-P$, by assigning a velocity $dL_{y}/dt=k_{p}(P_{t}-\overline{P})$ to the box edges with normals $\pm\mathbf{e}_{y}$, where $\overline{P}$} is the average pressure calculated between the control inputs, $P_{t}$ is the target vertical pressure, and $k_{p}$ is the proportional control gain. We set the target pressure so that the contact stiffness number $\kappa=\min(\alpha,1)(E/P)^{2/3}>1.1\cdot 10^{3}$, where $E$ is the Young modulus of the grains. Compared to [19] we scale the contact stiffness number by the aspect ratio, in order to guarantee a small overlap even for oblate particles. Doing so ensures we are in the hard grain limit [19, 50] and the relevant dimensionless number for several rheological properties [51] is the inertial number $I=\dot{\gamma}d_{eq}\sqrt{\rho_{p}/P}$, where $\dot{\gamma}$ is the shear rate, $\rho_{p}$ is the particle density, $d_{eq}=2a\alpha^{1/3}$ is the equivalent volume sphere diameter. Nevertheless, finite particle elasticity affects both coordination and volume fraction [52, 51, 29, 53]. We set the Poisson’s ratio $\nu=0.3$, which sets the ratio of tangential to normal stiffness $k_{t}/k_{n}\approx 0.41$, see Appendix A for more detail.
@@
$N=2000$ particles are inserted in the box, with approximately $12$ particles per dimension and, with a $20\%$ degree of polydispersity for the value of the smallest semi-axis around the mean and none in the aspect ratio. All simulations start at an initial concentration $\phi_{0}=0.4$ with isotropic random distribution and orientation of grains, equal edges of the box $L_{x}=L_{y}=L_{z}$, and then the box is compressed in the $y$ direction to reach the desired pressure. The chosen number of particles is sufficient to characterize the rheology, see Appendix C. We set the coefficient of restitution $e=0.1$, insofar at the high packing fractions ($\phi>0.45$) we investigate, its influence in the range $[0.1,0.9]$ on several rheological quantities is minor [19, 50, 16, 54, 32, 51, 53].
@@ III.1 Alignment and Rotation
$\mathbf{Q}=\int_{\mathbb{S}^{2}}\left(\mathbf{p}\otimes\mathbf{p}-{\mathbf{I}\over 3}\right)\psi(\mathbf{p})d\mathbf{p}\,$, (3)
@@
where $\mathbf{I}$ is the identity matrix, and $\psi(\mathbf{p})$ is the probability distribution of $\mathbf{p}$ on the unit sphere, $\mathbb{S}^{2}$. The uniaxial nematic order parameter is defined as $S_{2}=3\lambda_{\text{max}}/2$, where $\lambda_{\text{max}}$ is the maximum eigenvalue of $\mathbf{Q}$. As shown in Figure 2a, $S_{2}$ increases from 0 for an isotropic distribution when $r_{g}\to 0$, to values close to $1$ for elongated and flattened particles, where a value $S_{2}=1$ means that all particles are aligned in the same direction. Additionally, we define as weakly anisotropic configurations where $0<S_{2}<0.35$.
All non-spherical particles exhibit some degree of nematic alignment. Furthermore, we verified that no persistent smectic phase exists by computing the transverse correlation function of the center of mass density field [7], whose rapid decay confirms the absence of long-range translational order.
@@
One might expect that microscopic friction broadens the distribution of particle orientations, thereby reducing nematic order. However, for particles with low to moderate eccentricity, the highest degree of nematic ordering is not observed in the frictionless limit, but at an intermediate friction coefficient of $\mu_{p}=0.1$. This underscores the dual role of friction: while it typically disrupts alignment in the nematic phase, it can promote alignment in the isotropic phase via torque-induced reorientation. Finally, notice how for small eccentricity, $r_{g}<0.1$ the order parameter is within the weakly anisotropic case. Interestingly, oblate particles align slightly better up to moderate values of friction coefficient $\mu_{p}\leq 0.4$, but worse at high friction values with respect to their reciprocal prolate ones.
@@
As reported in several studies [14, 8, 9, 27, 28, 20, 21, 22, 18, 10] grains do not simply align in the direction of least resistance to the flow, along the streamlines, but at a certain angle in the extension direction. Defining the mean angle of inclination of the major axis with respect to the flow direction projected in the $xy$ plane, $\theta_{x}$, 111for oblate ellipsoids since there are are two major axes we track the symmetry axis and then compute the corresponding angle for the major axis as $\pi/2-\theta_{x}$, we report data in Figure 2b. The higher $|r_{g}|$ the more the particles orient their major axis parallel to the flow direction. The effect of friction is also highly non-linear: $\theta_{x}$ is lower for $0.1<\mu_{p}<10$ at small degrees of eccentricity, but for high flattening and elongation, the minimum $\theta_{x}$ is found for high values of friction, when alignment is less intense. The existence of this alignment angle, known as the Leslie angle [56] in liquid crystal theory, is well-established. However, its microscopic origins in granular matter are not fully understood, and predicting its value from fundamental particle properties is an open problem. Recent work [57] has started to illuminate these origins, linking the angle’s stability to collisional noise in the shear flow.
@@
Another way to look at the emergence of alignment is to track the the average vorticity-component of angular velocity $\langle\omega_{z}\rangle$, see Figure 2c. When friction is present, spherical particles rotate on average as a fluid particle in shear flow [19] $\omega_{z}\approx\dot{\gamma}/2$. For non-spherical particles, elongation provides a geometric constraint, effectively hampering their rotation. For all values of friction, oblate particle rotate slightly less than prolate ones for the same degree of eccentricity. One possible explanation is that while some prolate particles orient their major axis along the vorticity direction, resulting in a log-rolling behavior, this is not observed for oblate particles (data not shown). This behavior is analogous to that seen in simple shear flow in a dilute viscous fluid, where rotation about the vorticity axis is stable for prolate spheroids when their minor axis aligns with it, but for oblate spheroids, stability occurs when their major axis is aligned instead [58]. In the high friction limit, yellow curve, the rotation approaches the one of the corresponding fluid particle, regardless of the shape. This is due to the lower packing fraction attained in this scenario, see Figure 7. To conclude, for non-spherical particles we notice that the average rotation is anticorrelated with the orientational order [7], but not directly related to the particles’ preferred orientation.
@@ III.2 Dissipation Regime diagram and fluctuations
As a function of the inertial number and the microscopic friction, the leading dissipative mechanism of granular flows and dense suspensions can be either due to normal inelasticity, as viscoelastic damping or to frictional sliding [32, 33, 22].
We can split the dissipated power accordingly in a normal component $\mathcal{P}_{n}$, and a tangential one due to frictional sliding $\mathcal{P}_{t}$, in formulas:
@@
where $F_{i}$ is the force due to contact $i$, $N_{c}$ is the total number of contacts, $u_{i}$ is the relative velocity of the grains at the contact point, $F^{n,\gamma}$ is the normal component of the force due to damping and $N_{s}$ is the number of sliding contacts, those which satisfy $F_{i}^{t}=\mu_{p}F_{i}^{n}$. Figure 3 shows for each aspect ratio the regime diagram of the dissipation as the level set curve where $\mathcal{P}_{t}=\mathcal{P}_{n}$. The parametric space has three regions: the normal dissipation regime for low $\mu_{p}$ and high $I$, the rolling or highly frictional regime for high $\mu_{p}$ and high $I$ and the frictional sliding regime for intermediate $\mu_{p}$ and low enough $I$. For $\alpha>1$ the diagram closely resembles the one obtained for sphero-cylinders in [22]. Here, we extend this diagram by also including $\alpha<1$. For values of $|M_{\alpha}|>1.2$ the crossover from the normal dissipation to the sliding regime occurs at higher inertial number the higher the absolute value of the shape ratio (see the inset). Regarding the boundary between the sliding and rolling regime, non-spherical particles start to roll at higher values of $I$ and $\mu_{p}$.
Figure 3: Dissipation regime diagram of granular shear flow as a level set curve where the frictional and normal dissipation are equal. Left lines mark the transition between the normal dissipation and the sliding frictional regime, while right ones the transition to the rolling regime, where virtually all contacts are sticking. The black dashed line has slope 2 as predicted by [32] for spheres. For all non-spherical particles the sliding regime is wider than for spherical ones, indicating non-spherical particles slide over a larger range of parameters. (Inset) Slice of $I_{c}$ at the normal dissipation-sliding transition for $\mu_{p}=0.01$, showing non monotonicity in $\alpha$. Slightly aspherical particles display a reduced normal dissipation regime.
To rationalize our findings we visualize the power dissipation normalized by the injected power into the system, given by $\dot{\gamma}\sigma_{xy}\Omega\sim\dot{\gamma}P\Omega$, in Figure 4. We observe that, up to moderate friction coefficients ($\mu_{p}\leq 0.4$), Figure 4a, the power dissipated due to inelasticity decreases as particles become more elongated or flattened for aspect ratios $M_{\alpha}>1.2$.
Owing to the system’s structural anisotropy, assuming all particles are perfectly aligned with their major axis along the flow direction, the cross-sectional areas of each grain are listed in Table 1.
@@
Oblate $\alpha$ 11 $\alpha$ $\alpha$ 11
Prolate 11 $\alpha$ $\alpha$ $1/\alpha$ $1/\alpha$
Among them, the gradient-vorticity cross-section $A_{yz}$ is the smallest, implying that collisions are least likely in the flow direction. Defining the particles’ velocity fluctuation as the excess over the linear velocity profile $\delta\bm{v}=\bm{v}-\dot{\gamma}y\mathbf{e}_{x}$, velocity fluctuations along the flow direction, $\delta v_{x}$, tend to persist longer, due to the reduced cross-sectional area. Moreover, as reported in [16] at high volume fractions and seen in Figure 5a and b, the dominant fluctuation is $\delta v_{x}$, indicating that most of the fluctuating kinetic energy is in the flow direction. Taken together, these effects offer a simple and consistent explanation for the reduced normal power dissipation observed in systems with anisotropic particles: lower collision frequency along the dominant fluctuation direction leads to less energy lost through collisions. This simple geometrical argument also clarifies why, for flattened particles, fluctuations in the vorticity direction are of the same magnitude as those in the flow direction, as their respective cross-sections are equal, $A_{xz}=A_{xy}$.
@@
Previous scaling arguments for circular particles estimated that the cross-over from normal dissipation to sliding regime occurs at $I_{c}\sim\mu_{p}^{2}$ [32], as indicated by the dashed line in Figure 3. In our case, we find roughly $I_{c}\sim f(\alpha)\mu_{p}^{2}$, where $f(\alpha)$ a non-monotonic function, with global minimum in $\alpha=1.0$. The reduced normal dissipation of aligned particles, shifts the transition to higher $I_{c}$ for both elongated [22] and flattened particles. For the latter, their slightly higher alignment shifts the transition at slightly higher inertial number compared to the respective reciprocal shapes.
@@
From Figure 3, we also observe that at a fixed inertial number $I$, the transition from rolling to sliding regime occurs at higher values of $\mu_{p}$ for nonspherical particles. Oblate shapes start rolling at higher $\mu_{p}$ than prolate ones with similar eccentricity, as they experience more anisotropic fabric and comparatively stronger forces (Sections III.5, III.6). Interestingly, this second transition from sliding to rolling is monotonic across all $M_{\alpha}$, indicating that in the rolling regime, lower volume fraction and less collective motion reduce geometric constraints; see Sections III.3 and III.4.
@@
Figure 4: (a) Normalized normal and (b) tangential dissipation for $I\approx 10^{-2}$ and different values of microscopic friction. On the left, data at $\mu_{p}<0.4$, on the right $\mu_{p}\geq 0.4$. In panel (a.1), the red (black) line represents the best-fit power law for frictionless prolate (oblate) particles, with a slope of 0.90-0.90, $(0.96)$. For particles with $\mu_{p}=0.1$, panel (b.1), the corresponding slopes are 0.30-0.30 for prolate and +0.33+0.33 for oblate particles. From (b.1, b.2), it is evident that the tangential dissipation increases up to a value of $\mu_{p}$ between $0.1$ and $1.0$, depending on the shape, and then reduces for higher values.
@@
Figure 5: (a) Ratio of the streamline-wise velocity fluctuation $\delta v_{x}$ to the gradient-wise velocity fluctuation $\delta v_{y}$, (b) and to spanwise vorticity $\delta v_{z}$. (c) Ratio of rotational to translational kinetic energy, where the dashed line indicates a ratio of one. Different microscopic friction coefficients, with the same color legend as in Figure 4. For low values of friction the flow component of the fluctuation dominates for prolate particles, whereas for oblate ones, both the flow and the vorticity ones are dominant. In the sliding regime the fluctuations in the vorticity direction are reduced, as friction stabilizes that translation. For low values of friction and eccentricity the rotational fluctuations dominate. At high values of friction energy tends to be equally distributed between translation and rotation, as well as between the translational degrees of freedom. Data taken at $I\approx 0.01$.
@@
Figure 6: (a.1-6) Intensity of normalized velocity fluctuations, $\delta v=|\delta\bm{v}|$, and (b.1-6) angular velocity fluctuations, $\delta\omega=|\delta\bm{\omega}|$. Results are shown for $\mu_{p}=0$ (left), $0.4$ (center), and $10.0$ (right). Black lines indicate best-fit power laws for spherical particles, with slopes of 0.49-0.49, 0.19-0.19, and 0.46-0.46 (a.1-6), and 0.54-0.54, 0.25-0.25, and 0.47-0.47 (b.1-6), where in the frictionless case the fit is performed on $\alpha=1.5$. Splitting the data for elongated and flattened particles makes the symmetry of reciprocal shapes less evident, but clearly shows the monotonic trend.
[…]
@@
In Figure 6 we report the intensity of velocity fluctuations in the different regimes. Defining $\mathcal{L}=\delta v/\dot{\gamma}d_{eq}$, the so-called lever effect, which measures the dimensionless intensity of the velocity fluctuations ($\delta v$), in the normal dissipation and rolling regimes, $\mathcal{L}\sim I^{-\xi}$, with $\xi\approx 0.5$ as reported by multiple studies [50, 32, 33, 61, 62]. In the sliding regime however, we find $\mathcal{L}\sim I^{-0.19}$ for spherical particles, slightly lower than the one reported in 2D [32]. For elongated particles, the fluctuations are reduced even more in the sliding regime, confirming the more coherent flow they undergo. In the frictional-rolling regime the shape of the particles has very little influence on the velocity fluctuations, signaling that all particles behave in a similar way as alignment is reduced, and the fluctuations tend to become isotropic, recall Figure 5.